# The environment of TeV halo progenitors

### v1

@@ Abstract
Context. TeV halos are extended very-high-energy gamma-ray sources found around some middle-aged pulsars. The emission spanning several tens of parsecs suggests an efficient confinement of the ultra-relativistic lepton pairs produced by pulsars in their vicinity. The physical mechanism responsible for this suppressed transport has not yet been identified. In some scenarios, pair confinement may be linked to the medium the pulsars are located in.
Aims. We aim at understanding which type of medium pulsars are probing over their lifetime.
Methods. We developed a model for the environment probed by moving pulsars, from their birth in core-collapse explosions – where they receive a natal kick – until their entry into the interstellar medium. The model involves: (i) a Monte-Carlo sampling of the properties of the massive-star progenitors of pulsars; (ii) a calculation of the structure of the surrounding medium shaped by these progenitors, for the two cases of isolated stars and star clusters; (iii) a computation of the evolution of supernova remnants in these parent environments. Ultimately, from a distribution of neutron star kick velocities, we assess in which medium pulsars are located as a function of time. We first derive the statistical properties of a fully synthetic Galactic population, and then apply the model to a selection of known pulsars to assess the likely nature of their environment.
Results. We show that pulsars escape into the ISM at around $200\leavevmode\nobreak\ \mathrm{kyr}$, significantly later than the values most commonly used in the literature. The majority of known pulsars with a confirmed TeV halo have high probabilities of still being in their parent environment, which suggests that efficient pair confinement is connected to the region influenced by progenitor stars. In order to test this idea, we provide the probability of still residing in the parent environment for a list of known pulsars.
@@ 1 Introduction
Since the discovery of TeV halos around the Geminga and B0656+14 pulsars by the HAWC experiment in 2017 (Abeysekara et al. 2017), and around J0622+3749 by LHAASO in 2021 (Aharonian et al. 2021), theoretical efforts have been dedicated to understanding this source class. These very high-energy ($\gtrsim 100\leavevmode\nobreak\ \mathrm{GeV}$) gamma-ray emissions are produced by the inverse-Compton scattering of leptons accelerated in the pulsar wind nebula of the pulsar with the cosmic microwave background and the interstellar photon fields (Di Mauro et al. 2019). Known halos originate from middle-aged pulsars ($\sim 100\leavevmode\nobreak\ \mathrm{kyr}$) that are expected to have escaped their parent supernova remnant (SNR) into the interstellar medium (ISM) (Abeysekara et al. 2017; Fang 2022; Amato & Recchia 2024). Surprisingly, the gamma-ray emission hints at a stronger confinement – with a suppression factor $\gtrsim 100$ with respect to average conditions in the Galaxy – of the high energy ($\gtrsim 100\leavevmode\nobreak\ \mathrm{TeV}$) particles around a region of radius $20-30\leavevmode\nobreak\ \mathrm{pc}$ around the pulsars.
@@
The debate regarding how many TeV halos should be detectable has remained active since their discovery. Several studies expected the number of detected pulsar halos to grow rapidly (Linden et al. 2017; Linden & Buckman 2018). However the number of candidate TeV halos grew to just a dozen after almost 10 years of searches, with most of them needing further spectral and morphological studies to be confirmed. A high occurrence rate of TeV halos around middle-aged pulsars also seems disfavoured by the local positron flux and the properties of the known population of nearby pulsars (Martin et al. 2022).
@@
This discovery has consequences not only in gamma-ray astronomy but also in cosmic-ray physics. For example, before 2017, there was an effort made to understand whether pulsars or dark matter could explain the anomalous positron excess in cosmic rays (Yuan et al. 2015; Boudaud et al. 2015). The detection of TeV halos led to the idea that the large-scale cosmic-ray diffusion coefficient of the Galaxy – deduced from secondary/primary cosmic-ray nuclei fluxes and mainly probing the large diffusive halo of the Milky Way – is not representative of transport conditions in the disc. As a consequence of this lower diffusion coefficient, cosmic-ray leptons would take more time diffuse between nearby sources and the Earth, suffer stronger energy losses especially at the highest energies, which would make nearby pulsars unable to explain the positron excess (Abeysekara et al. 2017). Later, more refined studies showed that pulsars could still explain the positron excess if propagation around them is described in a two-zone framework, with suppressed diffusion within tens of parsec of the source – the halo – and normal standard diffusion at larger distances  (Tang & Piran 2019; Schroer et al. 2023).
@@
The theoretical efforts to explain the suppressed diffusion can be divided into three categories, depending on whether the suppression is actually necessary, or whether it is considered necessary but either linked to the pulsar progenitor or not. To the first category belong, for example, models of anisotropic transport along the magnetic fields surrounding the pulsar (Liu et al. 2019; De La Torre Luque et al. 2022), or models that take into account non purely diffusive transport (Recchia et al. 2021). To the second category belong, for example, models where the suppressed diffusion is attributed to turbulence self-generated by the lepton pairs emitted by the pulsar (Evoli et al. 2018; Mukhopadhyay & Linden 2022) through the excitation of plasma instabilities. To the third category – the one relevant for the present work – belong, for instance, the work by Fang et al. (2019), which proposed that the pulsars showing a TeV halo are still inside their parent SNRs, whose interior could be turbulent, or the work by Schroer et al. (2022), which proposed that the turbulence responsible for the formation of the TeV halo may be self-generated by protons accelerated by the SNR and escaping the system.
@@
The lepton pairs creating TeV halos are also expected to emit in other wavelengths, such as radio or X-rays through synchrotron emissions. Both should in principle allow to gain new information on the environment of the pulsar, especially the magnetic field strength or structure, and the emitting electron and positron population (Khokhriakova et al. 2024). In the case of Geminga, no X-ray synchrotron halo could be detected and  Manconi et al. (2024) gave an upper limit on the magnetic field strength around the pulsar.
@@
Despite intense interest in TeV halos, there remains no clear consensus on the environments that host TeV-halo pulsars– critically limiting our ability to pinpoint the physical mechanisms behind the observed suppression of cosmic-ray diffusion. This uncertainty is exemplified by Geminga, whose putative supernova remnant remains undetected, and by PSR B0656+14, which appears projected within the Monogem ring, itself a candidate parent SNR. Since the nature and level of turbulence–and thus particle propagation–can vary dramatically depending on the local environment, resolving this ambiguity is essential. In this work, we take a decisive step forward: by reconstructing the past trajectory and environment of each pulsar, we directly investigate the medium traversed throughout its life, providing crucial insights into the origin and properties of TeV halos.
It is commonly assumed that pulsars escape their parent SNR at about $60\leavevmode\nobreak\ \mathrm{kyr}$ due to their natal kick velocity (Amato & Recchia 2024), based on the evolution of a single SNR evolving in the Sedov-Taylor stage in a warm neutral ISM. This leads to the hypothesis that the middle-aged pulsars creating the halos are all in the ISM.
@@
However, pulsars are born from massive stars that, whether isolated or in groups, blow powerful winds during their lives. The stellar winds carve rarefied, hot and turbulent bubbles that can extend up to several hundreds of parsecs (Weaver et al. 1977; Mac Low & McCray 1988; Vieu et al. 2022). Then, the supernova explosion following the birth of the pulsar ejects stellar material inside the bubble which further modifies the surrounding medium. Accounting for this peculiar environment challenges the usual understanding of the environment of pulsars. Stellar bubbles and their interaction with the SNR were never studied in the literature relative to TeV halos.
@@
In this paper, we present in Sect. 2 a model for the environment probed by pulsars following their formation and the impulse they receive in core-collapse supernova explosions of massive-star progenitors. We provide in Sect. 3 the statistical description of the environment of pulsars as a function of time, for a fully synthetic Galactic population. In Sect. 4 we apply our model to a selection of known pulsars to determine statistically the nature of the medium they are most likely to reside in, and we discuss in more detail a number of individual objects including the current sample of established TeV halos. Finally, we conclude in Sect. 5.
@@ 2 Model
We here introduce the assumptions made about the pulsar initial kick velocity and about the development of the surrounding medium in relation to the massive-star progenitors.
@@ 2.1 Pulsar kinematics
Pulsar kick velocities result from asymmetries in the SN explosion happening at the end of the life of the progenitor massive star (Lambiase & Poddar 2024). Assuming a uniform and rectilinear trajectory, the pulsar position depends on kick velocity $v_{\mathrm{kick}}$ and true age $t_{\mathrm{age}}$ as $r_{\mathrm{PSR}}=v_{\mathrm{kick}}t_{\mathrm{age}}$.
A recent work using parallax and proper-motion measurements of radio pulsar by Igoshev (2020) suggests using single and bimodal Maxwellian distributions for the kick velocities of old and young pulsars, respectively (young meaning a spin-down characteristic age $\tau_{\mathrm{c}}<3\leavevmode\nobreak\ \mathrm{Myr}$). Since the middle-aged pulsars that we consider in the present work have a characteristic age comprised in the range of $100-300\leavevmode\nobreak\ \mathrm{kyr}$, we adopted a bimodal Maxwellian distribution $f_{2\mathcal{M}}$:
@@
where $v$ is short for $v_{\mathrm{kick}}$, $\mathcal{M}(v)$ is the Maxwellian distribution, and the optimal parameters are $w=0.2$, $\sigma_{1}=128\leavevmode\nobreak\ \mathrm{km/s}$, and $\sigma_{2}=336\leavevmode\nobreak\ \mathrm{km/s}$ [^1] It is puzzling that a dip in the velocity distribution for young pulsars, at roughly $200\leavevmode\nobreak\ \mathrm{km/s}$, lies exactly where a peak is found in the velocity distribution for all pulsars. Disberg et al. (2025) find it more likely that the low-velocity peak in the bimodal velocity distribution is due to Poisson noise on the nearby pulsars instead of having a physical origin..
We assessed the impact of using different kick velocity distributions on a subset of our results by considering the total sample of pulsars in Igoshev (2020) or previous distributions adopted in the literature such as those from Faucher-Giguère & Kaspi (2006) and Hobbs et al. (2005).
@@ 2.2 Progenitor stars
We assumed that pulsars are born from massive stars of zero-age main-sequence (ZAMS) masses $M_{\mathrm{ZAMS}}$ between $8\leavevmode\nobreak\ \mathrm{M}_{\odot}<M_{\mathrm{ZAMS}}<40\leavevmode\nobreak\ \mathrm{M}_{\odot}$ (Sukhbold et al. 2016). Stars in this mass range follow a power-law initial-mass function (IMF) of index $-2.35$ (Kroupa & Weidner 2003). Stars with ZAMS masses in the interval $2\leavevmode\nobreak\ \mathrm{M}_{\odot}<M_{\mathrm{ZAMS}}<16\leavevmode\nobreak\ \mathrm{M}_{\odot}$ are considered B stars, while stars with higher masses are O stars. After reaching the end of the main sequence (MS) phase, massive stars enter the red supergiant (RSG) and Wolf-Rayet (WR) phases. As progenitor properties, we use the results of Seo et al. (2018) and references therein, which are based on stellar evolution calculations by Ekström et al. (2012) for non-rotating stars.
The MS and RSG phase durations $\tau_{\mathrm{MS}}$ and $\tau_{\mathrm{RSG}}$ depend on $M_{\mathrm{ZAMS}}$ and are given in Myr as (Zakhozhay 2013; Seo et al. 2018):
[…]
$+0.19\left(\log M_{\mathrm{ZAMS}}\right)^{3}-0.057\left(\log M_{\mathrm{ZAMS}}\right)^{4}.$ (2)
@@
The fraction of the lifetime spent as RSG decreases strongly with the mass and amounts to 17% for a $8\leavevmode\nobreak\ \mathrm{M}_{\odot}$ B star or 2% for a $40\leavevmode\nobreak\ \mathrm{M}_{\odot}$ O star.
Non-rotating stars with masses $\geq 25\leavevmode\nobreak\ \mathrm{M}_{\odot}$ also experience a WR phase that lasts a few hundreds of thousands of years, with continuous emission of even more powerful winds. We did not consider the WR phase in our model for the following reasons. First, given the stellar mass range and initial mass function considered, only a very small fraction of massive stars will actually experience a WR phase. Second, WR stars in the mass range of interest lose a very small amount of mass and during a very short time, so the impact on the global dynamics of the environment will not be noticeable since the WR winds will mainly propagate within the RSG winds and not affect so much the global structure of the surrounding medium. Hence, we neglected the WR phase in our calculations.
The total mass lost $\Delta M$ during each phase is given in $\mathrm{M}_{\odot}$ by:
@@
For the instantaneous wind mass-loss rate $\dot{M}$ (in $\mathrm{M}_{\odot}/\mathrm{yr}$) and velocity $u_{\mathrm{w}}$ (in $\mathrm{km}/\mathrm{s}$) in each phase, we followed the prescription of Seo et al. (2018):
@@
$u_{\mathrm{w,RSG}}$ $\approx 1.9\cdot M_{\mathrm{RSG}}^{0.85}\leavevmode\nobreak\ \mathrm{km/s}.$ (10)
where the quantities are given as a function of the mass of the start at the beginning of each phase, so $M_{\mathrm{MS}}=M_{\mathrm{ZAMS}}$ and $M_{\mathrm{RSG}}=M_{\mathrm{ZAMS}}-\Delta M_{\mathrm{MS}}$. The wind luminosity $L_{\mathrm{w}}$ in $\mathrm{erg/s}$ derives from the above quantities as $L_{\mathrm{w}}=\dot{M}u_{\mathrm{w}}^{2}/2$, which yields for the MS:

### v2

@@ Abstract
Context. TeV haloes are extended sources of very-high-energy gamma rays found around some middle-aged pulsars. The emission spanning several tens of parsecs suggests an efficient confinement of the ultra-relativistic lepton pairs produced by pulsars in their vicinity. The physical mechanism responsible for this suppressed transport has not yet been identified. In some scenarios, pair confinement may be linked to the medium the pulsars are located in.
Aims. We aim at understanding the type of medium pulsars probe over their lifetime.
Methods. We developed a model for the environment probed by moving pulsars, from their birth in core-collapse explosions – where they receive a natal kick – until their entry into the interstellar medium. The model involves: (i) a Monte-Carlo sampling of the properties of the massive-star progenitors of pulsars; (ii) a calculation of the structure of the surrounding medium shaped by these progenitors for the two cases of isolated stars and star clusters; and (iii) a computation of the evolution of supernova remnants in these parent environments. Ultimately, from a distribution of neutron star kick velocities, we assess the medium in which pulsars are located as a function of time. We first derived the statistical properties of a fully synthetic Galactic population and then applied the model to a selection of known pulsars to assess the likely nature of their environment.
Results. We show that pulsars escape into the interstellar medium at around $300\penalty 10000\ \mathrm{kyr}$, significantly later than assumed in the literature. Given our assumptions, all known pulsars with a confirmed TeV halo have high probabilities of still being in their parent environment, which suggests that efficient pair confinement is connected to the region influenced by progenitor stars. To test this, we provide the probability that known pulsars still reside in their parent environment for a list of known pulsars.
@@ 1 Introduction
Since the discovery of TeV haloes around the Geminga and B0656+14 pulsars by the HAWC experiment in 2017 (Abeysekara et al. 2017), and around J0622+3749 by LHAASO in 2021 (Aharonian et al. 2021), theoretical efforts have been dedicated to understanding this source class. These very high-energy ($\gtrsim 100\penalty 10000\ \mathrm{GeV}$) gamma-ray emissions are produced by the inverse-Compton scattering of leptons accelerated in the pulsar wind nebula of the pulsar with the cosmic microwave background and the interstellar photon fields (Di Mauro et al. 2019). Known haloes originate from middle-aged pulsars ($\sim 100\penalty 10000\ \mathrm{kyr}$) expected to have escaped their parent supernova remnant (SNR) into the interstellar medium (ISM) (Abeysekara et al. 2017; Fang 2022; Amato and Recchia 2024). Surprisingly, the gamma-ray emission hints at a stronger confinement – with a suppression factor $\gtrsim 100$ with respect to average conditions in the Galaxy – of the high energy ($\gtrsim 100\penalty 10000\ \mathrm{TeV}$) particles around a region of radius $20-30\penalty 10000\ \mathrm{pc}$ around the pulsars.
@@
The debate regarding the number of detectable TeV haloes has remained active since their discovery. Several studies expected the number of detected pulsar haloes to grow rapidly (Linden et al. 2017; Linden and Buckman 2018). However, after almost 10 years of searches, the number of candidate TeV haloes grew to just a dozen with most requiring further spectral and morphological studies for confirmation. A high occurrence rate of TeV haloes around middle-aged pulsars also seems disfavoured by the local positron flux and the properties of the known population of nearby pulsars (Martin et al. 2022).
@@
This discovery has consequences not only in gamma-ray astronomy but also in cosmic-ray physics. For example, before 2017, an effort was made to understand whether pulsars or dark matter could explain the anomalous positron excess in cosmic rays (Yuan et al. 2015; Boudaud et al. 2015). The detection of TeV haloes led to the idea that the large-scale cosmic-ray diffusion coefficient of the Galaxy – deduced from secondary and primary cosmic-ray nuclei fluxes, mainly probing the large diffusive halo of the Milky Way – is not representative of transport conditions in the disc. As a consequence of this lower diffusion coefficient, cosmic-ray leptons would take more time to diffuse between nearby sources and the Earth, suffer stronger energy losses particularly at the highest energies, which would make nearby pulsars unable to explain the positron excess (Abeysekara et al. 2017). Later, more refined studies showed that pulsars could still explain the positron excess if propagation around them is described in a two-zone framework, with suppressed diffusion within tens of parsec of the source – the halo – and normal standard diffusion at larger distances (Tang and Piran 2019; Schroer et al. 2023).
@@
The theoretical efforts to explain the suppressed diffusion can be divided into three categories, depending on whether the suppression is actually necessary or considered necessary but either linked to the pulsar progenitor or not. Belonging to the first category, for example, are models of anisotropic transport along the magnetic fields surrounding the pulsar (Liu et al. 2019b; De La Torre Luque et al. 2022), or models that take into account non purely diffusive transport (Recchia et al. 2021). Belonging to the second category, are models where the suppressed diffusion is attributed to turbulence self-generated by the lepton pairs emitted by the pulsar (Evoli et al. 2018; Mukhopadhyay and Linden 2022) through the excitation of plasma instabilities. Belonging to the third category – the one relevant for the present work – include, for instance, the work by Fang et al. (2019), which proposes that pulsars exhibiting a TeV halo remain inside their parent SNRs, whose interiors could be turbulent, and the work by Schroer et al. (2022), which suggests that the turbulence responsible for the formation of the TeV halo may be self-generated by protons accelerated by the SNR and escaping the system.
@@
The lepton pairs creating TeV haloes are also expected to emit in other wavelengths, such as radio or X-ray through synchrotron emission. Both should in principle allow us to gain new information on the environment of the pulsar, especially its magnetic field strength or structure (Liu et al. 2019a), and the emitting electron and positron population (Khokhriakova et al. 2024). In the case of Geminga, no X-ray synchrotron halo could be detected, and Manconi et al. (2024) gave an upper limit on the magnetic field strength around the pulsar.
@@
Despite intense interest in TeV haloes, there remains no clear consensus on the environments that host TeV-halo pulsars – critically limiting our ability to pinpoint the physical mechanisms behind the observed suppression of cosmic-ray diffusion. This uncertainty is exemplified by Geminga, whose putative SNR remains undetected, and by PSR B0656+14, which appears projected within the Monogem ring, itself a candidate parent SNR. Since the nature and level of turbulence – and thus particle propagation – can vary dramatically depending on the local environment, resolving this ambiguity is essential. In this work, we take a decisive step forward: by reconstructing the past trajectory and environment of each pulsar, we directly investigate the medium traversed throughout its life, providing crucial insights into the origin and properties of TeV haloes.
van der Swaluw et al. (2003) computed the escape time $t_{\mathrm{esc}}$ at which a pulsar with speed $v_{\mathrm{kick}}$ reaches the ISM when escaping from a SNR of energy $E_{\mathrm{SN}}$ propagating in the uniform ISM of density $n_{\mathrm{ISM}}$ and found
$t_{\mathrm{esc}}=14\left(\frac{E_{\mathrm{SN}}}{10^{51}\penalty 10000\ \mathrm{erg}}\right)^{1/3}\left(\frac{v_{\mathrm{kick}}}{1000\penalty 10000\ \mathrm{km/s}}\right)^{-5/3}\left(\frac{n_{\mathrm{ISM}}}{1\penalty 10000\ \mathrm{cm}^{-3}}\right)^{-1/3}\penalty 10000\ \mathrm{kyr}.$ (1)
This leads to the hypothesis that the middle-aged pulsars creating the haloes are all in the ISM.
@@
However, pulsars are born from massive stars that, whether isolated or in groups, blow powerful winds during their lives. The stellar winds carve rarefied, hot, and turbulent bubbles that can extend up to several hundreds of parsecs (Weaver et al. 1977; Mac Low and McCray 1988; Vieu et al. 2022). Then, the supernova explosion following the birth of the pulsar ejects stellar material inside the bubble, which further modifies the surrounding medium. Accounting for this peculiar environment challenges the usual understanding of the environment of pulsars. Stellar bubbles and their interaction with the SNR were never studied in the literature relative to TeV haloes.
@@
In this paper, we present in Sect. 2 a model for the environment probed by pulsars following their formation and the impulse they receive in core-collapse supernova explosions of massive-star progenitors. We provide in Sect. 3 the statistical description of the environment of pulsars as a function of time, for a fully synthetic Galactic population. In Sect. 4 we apply our model to a selection of known pulsars to statistically determine the nature of the medium they are most likely to reside in, and we discuss in more detail a number of individual objects including the current sample of established TeV haloes. Finally, we conclude in Sect. 5.
@@ 2.1 Pulsar kinematics
Pulsar kick velocities result from asymmetries in the SN explosion happening at the end of the life of the progenitor massive star (Lambiase and Poddar 2024). Assuming a uniform and rectilinear trajectory, the pulsar position depends on kick velocity $v_{\mathrm{kick}}$ and true age $t_{\mathrm{age}}$ as $r_{\mathrm{PSR}}=v_{\mathrm{kick}}t_{\mathrm{age}}$.
A recent work using parallax and proper-motion measurements of radio pulsar by Igoshev (2020) suggests using single and bimodal Maxwellian distributions for the kick velocities of old and young pulsars, respectively (young meaning a spin-down characteristic age $\tau_{\mathrm{c}}<3\penalty 10000\ \mathrm{Myr}$). Since the middle-aged pulsars that we consider in the present work have a characteristic age comprised in the range of $100-300\penalty 10000\ \mathrm{kyr}$, we adopted a bimodal Maxwellian distribution $f_{2\mathcal{M}}$:
@@
where $v$ is short for $v_{\mathrm{kick}}$, $\mathcal{M}(v)$ is the Maxwellian distribution, and the optimal parameters are $w=0.2$, $\sigma_{1}=56\penalty 10000\ \mathrm{km/s}$, and $\sigma_{2}=336\penalty 10000\ \mathrm{km/s}$ [^1] It is puzzling that a dip in the velocity distribution for young pulsars, at roughly $200\penalty 10000\ \mathrm{km/s}$, lies exactly where a peak is found in the velocity distribution for all pulsars. Disberg et al. (2025) find it more likely that the low-velocity peak in the bimodal velocity distribution is due to Poisson noise on the nearby pulsars instead of having a physical origin..
We assessed the impact of using different kick velocity distributions on a subset of our results by considering the total sample of pulsars in Igoshev (2020) or previous distributions, such as those from Faucher-Giguère and Kaspi (2006) and Hobbs et al. (2005), adopted in the literature.
@@ 2.2 Progenitor stars
We assumed that pulsars are born from massive stars of zero-age main-sequence (ZAMS) masses $M_{\mathrm{ZAMS}}$ between $8\penalty 10000\ \mathrm{M}_{\odot}<M_{\mathrm{ZAMS}}<20\penalty 10000\ \mathrm{M}_{\odot}$ (Sukhbold et al. 2016). Stars in this mass range follow a power-law initial-mass function (IMF) of index 2.35-2.35 (Kroupa and Weidner 2003). Stars with ZAMS masses in the interval $2\penalty 10000\ \mathrm{M}_{\odot}<M_{\mathrm{ZAMS}}<16\penalty 10000\ \mathrm{M}_{\odot}$ are considered B stars, while stars with higher masses are O stars. After reaching the end of the main-sequence (MS) phase, massive stars enter the red supergiant (RSG) and Wolf-Rayet (WR) phases. As progenitor properties, we used the results of Seo et al. (2018) and references therein, which are based on stellar evolution calculations by Ekström et al. (2012) for non-rotating stars.
The MS and RSG phase durations, $\tau_{\mathrm{MS}}$ and $\tau_{\mathrm{RSG}}$, depend on $M_{\mathrm{ZAMS}}$ and are given in megayears as (Zakhozhay 2013; Seo et al. 2018)
$\log\tau_{\mathrm{MS}}$ $\approx 9.96\,-\,3.32\log M_{\mathrm{ZAMS}}+0.63\left(\log M_{\mathrm{ZAMS}}\right)^{2}$
$+0.19\left(\log M_{\mathrm{ZAMS}}\right)^{3}-0.057\left(\log M_{\mathrm{ZAMS}}\right)^{.},$ (3)
@@
The fraction of the lifetime spent as RSG decreases strongly with the mass and amounts to 17% for a $8\penalty 10000\ \mathrm{M}_{\odot}$ B star or 7% for a $20\penalty 10000\ \mathrm{M}_{\odot}$ O star.
We did not account for the WR phase as it only affects stars with masses $\geq 25\penalty 10000\ \mathrm{M}_{\odot}$.
The total mass lost $\Delta M$ during each phase is given in $\mathrm{M}_{\odot}$ by
@@
For the instantaneous wind mass-loss rate $\dot{M}$ (in solar masses per year) and velocity $u_{\mathrm{w}}$ (in kilometres per second) in each phase, we followed the prescription of Seo et al. (2018):
@@
$u_{\mathrm{w,RSG}}$ $\approx 1.9\cdot M_{\mathrm{RSG}}^{0.85}\penalty 10000\ \mathrm{km/s},$ (11)
where the quantities are given as a function of the mass of the start at the beginning of each phase, so $M_{\mathrm{MS}}=M_{\mathrm{ZAMS}}$ and $M_{\mathrm{RSG}}=M_{\mathrm{ZAMS}}-\Delta M_{\mathrm{MS}}$. The wind luminosity $L_{\mathrm{w}}$ in ergs per second derives from the above quantities as $L_{\mathrm{w}}=\dot{M}u_{\mathrm{w}}^{2}/2$, which yields the following for the MS: